OpenStudy (anonymous):

Write the equation of the line that passes through (3, 6) and (0, –3) in slope-intercept form

OpenStudy (anonymous):

let me try that again from 0 to 3 is three steps right, from -3 to 6 is nine steps up, so slope is 9/3 = 3

OpenStudy (anonymous):

y - intercept is given to you as -3 so equation is \[y=3x-3\]

OpenStudy (anonymous):

actually the slope is 3

OpenStudy (anonymous):

Write the equation of the line that passes through (1, 3) and (4, 4) in standard form

OpenStudy (anonymous):

(3, 6) and (0, –3) m = 9/ 3 = 3 -3 = 0 + b => y = 3x - 3

OpenStudy (anonymous):

I remember everybody already solve your question about standard form.